Compute the extension in length of a wire 530 cm long and a diameter of 0.25 cm. When the body of mass 540 gm is applied vertically. The young modulus of given wire is 5x109 Dyne/cm2.


1
Expert's answer
2021-06-22T09:47:54-0400

The area is


"A=\frac{\pi d^2}{4}."

5·109 Dyne/cm2 is 5·108 Pa.

Find the stress:


"\sigma=\frac FA=\frac{mg}{A}."

The strain:


"\epsilon=\frac{\Delta l}{l}\rightarrow\Delta l=\epsilon l."

And the Young's modulus is


"E=\frac\sigma\epsilon\rightarrow\epsilon=\frac\sigma E\rightarrow\Delta l=\frac\sigma El,\\\space\\
\Delta l=\frac{4mg}{\pi d^2E}l=1.14\text{ cm}."